# Large Language Model Psychometrics: A Systematic Review of Evaluation, Validation, and Enhancement

### v1

@@ 6.2 Data and Task Sources
Table 6: Overview of data and task sources: established inventories, custom-curated items, and synthetic items. For each data source, we list some examples for each construct.
@@
Established psychometric inventories present well-validated and widely recognized tools. Representative inventories in LLM psychometrics include BFI and HEXACO for personality traits [Serapio-García et al., 2023, Barua et al., 2024]; SVS, PVQ, WVS, and VSM for values [Fischer et al., 2023, Miotto et al., 2022]; MFT and DIT for morality [Abdulhai et al., 2024, Khandelwal et al., 2024]; ANES and PCT for political attitudes and opinions [Argyle et al., 2023, Bernardelle et al., 2024]; cognitive reflection tests (CRT) for heuristics and biases [Hagendorff et al., 2023, Binz and Schulz, 2023]; False-Belief Tasks for ToM [Ullman, 2023, Chen et al., 2024d]; and many more. Established inventories provide a straightforward and standardized approach to psychometric evaluation. However, many studies question the reliability and validity of established inventories when used with LLMs, citing common issues of data contamination and response bias [Ye et al., 2025a]. Additionally, these simplified, structured tests may not adequately capture the complex task scenarios encountered by general-purpose AI.
@@
Human-authored and custom-curated items provide tailored psychometric tests that are often more relevant and applicable to LLMs. These items enable researchers to explore novel and unique dimensions of LLM capabilities more robustly. Shen et al. [2024] argue that SVS and PVQ are inadequate for measuring LLM value alignment because they lack supplementary values from alignment literature and contextual nuances of human-AI alignment. To address this, they created 11 AI-informed value statements, similar in style to SVS and PVQ, based on a systematic review of human-AI alignment literature. Similarly, Ceron et al. [2024] suggest that political questionnaires may yield unreliable or inconsistent results. They propose a series of tests to assess the reliability and consistency of LLMs’ stances on political statements, using a dataset of voting-advice questionnaires from seven EU countries annotated for policy issues. Regarding the evaluation of LLM morality, Hendrycks et al. [2021] curate a large-scale dataset of over 130,000 open-world scenarios requiring LLMs to make moral judgments. Another example is the ToM dataset based on naturally occurring spoken dialogues [Soubki et al., 2024], which is motivated by the potential misalignment between synthetic ToM benchmarks and actual human behavior. Despite their high ecological validity, custom-curated items require significant effort to develop and validate, which may limit their scalability and diversity.
@@
Synthetic items generated by AI models, mostly LLMs themselves, represent an emerging paradigm in LLM psychometrics. This method leverages the generative capabilities of LLMs to produce large-scale, diverse, and contextually rich items. Using synthetic items entails careful prompt engineering and additional validation.
Some studies prompt LLMs to modify or expand established inventories for better ecological validity, mitigating data contamination, and enabling more scalable and diverse testing. For example, Ren et al. [2024] rewrite self-report items into advice-seeking queries that better align with real-world human-AI interactions. Bhandari et al. [2025b] modify the established items into semantically equivalent ones to address the issue of data contamination. They validate the rewritten items by setting a threshold for semantic embedding similarity. Zhu et al. [2024a] transform the cognitive test items into more diverse ones, allowing for multifaceted analysis of LLMs’ cognitive abilities.
Other studies generate synthetic tests from scratch. Hadar-Shoval et al. [2025] validate their synthetic cognitive tests by showing a strong human performance correlation between LLM-generated tests and established tests. Ye et al. [2025a] generate value-eliciting prompts for LLM value measurement, subsequently confirming the reliability and validity of the test results. Jiang et al. [2024a] propose generative evolving testing, where an LLM-based Item Generator learns to generate items with specified difficulty. The moral dilemmas presented by Chiu et al. [2025] are created using GPT-4 and employed to evaluate the moral judgment of LLMs. When measuring SI of LLMs with agentic simulations, Zhou et al. [2024b], Mou et al. [2024] prompt LLMs to generate various components of the social interaction.
@@ 6.3 Prompting Strategies
Role-playing prompts, also known as persona or profiling prompts, incorporate specific demographic information or personal attributes into the LLM context during test administration.
Some studies utilize role-playing prompts to generate multiple participants from a single LLM for statistical analysis. Serapio-García et al. [2023] incorporate persona instructions and descriptions before each item when assessing LLM personality. This approach yields diverse responses from a single LLM, facilitating the computation of the reliability and validity of the test results for each LLM. Ye et al. [2025b] use value-anchoring prompts to guide LLM values, thereby generating hundreds of LLM participants for value measurement. The statistical analysis of these measurement results leads to the development of an LLM-specific value system.
Research suggests that LLMs embody a superposition of diverse personalities, values, and perspectives [Kovač et al., 2023]. This has motivated investigations into the personality adaptability of LLMs. Utilizing role-playing prompts, Lu et al. [2023], Jiang et al. [2023], La Cava and Tagarelli [2024] demonstrate that LLMs can dynamically transition between different personality types. However, the degree of steerability varies among LLMs [La Cava and Tagarelli, 2024], and some personality dimensions are less steerable than others [Li et al., 2024d, Jiang et al., 2023]. Kovač et al. [2024], Rozen et al. [2024] employ role-playing prompts to examine the stability and consistency of LLM values. Similarly, Münker [2024], Simmons [2023], Wright et al. [2024] incorporate moral or political profiles into prompts to evaluate LLMs’ ability to mimic these perspectives and their susceptibility to biases linked to specific groups.
Many cognitive benchmarks, particularly those related to social interactions, assign personas and social roles to LLMs during agentic simulations [Zhou et al., 2024b, Mou et al., 2024, Huang et al., 2023b]. The successful completion of these benchmark tasks heavily relies on the LLM’s ability to comprehend and emulate the assigned persona and social role. Persona prompts are also shown to affect the social-cognitive reasoning of LLMs when other task variables are held constant [Tan et al., 2024].
The susceptibility to role-playing prompts presents a double-edged sword. On one hand, they enable the emulation of diverse perspectives and behaviors, which can be beneficial for applications such as role-playing conversations and social simulations. On the other hand, these capabilities may result in inconsistencies and instability in the LLM’s values and perspectives. It may compromise the reliability and validity of test results and pose challenges to robustly aligning LLMs with human values.
@@
Performance-enhancing prompts are crafted to augment LLMs in psychometric evaluations. Hagendorff [2023] suggest using such prompts whenever possible when evaluating machine psychology. The Chain of Thought (CoT) prompting [Wei et al., 2022] is leveraged to reduce biases [Hagendorff et al., 2023], enhance social intelligence [Shapira et al., 2023a], and boost cognitive capabilities [Coda-Forno et al., 2024]. Variants of CoT, such as Emotional CoT [Li et al., 2024e], specifically target improvements in the emotional intelligence of LLMs. Emotional prompts, as introduced by Li et al. [2023a], enhance the general cognitive abilities of LLMs by incorporating emotional stimuli. Few-shot prompting [Brown et al., 2020], another versatile prompting technique, has been shown to improve ToM performance in LLMs [Moghaddam and Honey, 2023].
Additional research explores diverse prompting strategies for specific constructs. For instance, Zhou et al. [2024a] advocate for MFT-guided reasoning prompts for better moral alignment. Strategies like those proposed by Echterhoff et al. [2024], Sumita et al. [2024] aim to self-debias LLMs. ToM-specific enhancement strategies include SymbolicToM [Sclar et al., 2023] and temporal decomposition-based prompting [Sarangi et al., 2025, Hou et al., 2024]. In contrast, Zhao et al. [2025a] develop a testing framework utilizing self-reflection prompts to examine both explicit and implicit social biases in LLMs, where explicit bias measurement serves as a reflection on implicit bias.
@@ 6.4.1 Closed-Ended Output and Scoring
The evaluation of LLM performance on structured tests can be categorized into two methodological paradigms: logit-based probabilistic analysis and closed-ended output scoring. In multiple-choice or Likert-scale evaluations, some studies retrieve and analyze the distribution of token-level logits—particularly the first generated token—to infer the model’s latent personalities and opinions [Pellert et al., 2024, Santurkar et al., 2023], analyze response entropy [Dominguez-Olmedo et al., 2024], and compute the distributional alignment between LLM outputs and human behavioral data [Arehalli et al., 2022]. Psycholinguistic tests especially rely on logit-based metrics. A notable example is surprisal, defined as the negative log probability of a token given its context. It quantifies the predictability or cognitive effort associated with language processing, enabling researchers to map LLM outputs to human-like uncertainty patterns [Wang et al., 2024b, Steuer et al., 2023].
@@ 6.4.2 Open-Ended Output and Scoring
Scoring open-ended outputs is more challenging. Related scoring schemes can be generally categorized into 1) rule-based scoring, 2) model-based scoring, and 3) human scoring.
@@
Rule-based scoring predominantly relies on the lexical hypothesis [Allport and Odbert, 1936], which posits that the meaning and relevance of responses can be ascertained by analyzing the presence and frequency of specific keywords or phrases within the text. Jiang et al. [2024b] utilize the LIWC (Linguistic Inquiry and Word Count) features [Pennebaker et al., 2001] to evaluate the personalities of LLMs. Similarly, Fischer et al. [2023] employ a theory-driven value dictionary [Ponizovskiy et al., 2020] to assess LLM values. Nevertheless, lexicon-based scoring has been demonstrated to be limited in capturing semantic nuances [Ye et al., 2025a]. Some other scoring rules extend beyond simple keyword matching. For example, Healey et al. [2024] automatically identify biases by analyzing whether responses deviate from equivalent treatment.
@@
Model-based scoring is prevalent in current unstructured testing due to its flexibility and scalability. Several studies have focused on training models specifically to evaluate LLM responses. For instance, Hilliard et al. [2024] fine-tuned BERT variants [Devlin et al., 2019] using the MyPersonality dataset [^5] https://sites.google.com/michalkosinski.com/mypersonality. to score LLM personalities. Similarly, Ye et al. [2025a], Yao et al. [2024, 2025b], Sorensen et al. [2024a] fine-tune LLMs with established psychometric inventories [Ren et al., 2024] and/or human annotations to classify the value valence of LLM responses. Scoring can occur at the item level, as exemplified by Generative Psychometrics, which consists of item parsing, scoring, and aggregation to evaluate LLM values [Ye et al., 2025a]. Alternatively, scoring can be conducted at the response level, assessing the LLM response in its entirety [Yao et al., 2024, 2025b]. Some models are trained to score according to specific theoretical frameworks, e.g., Schwartz’s Value Theory [Yao et al., 2024], while others are designed to be generalists, utilizing a broader range of fine-tuning data and leveraging the prior knowledge of LLMs [Ye et al., 2025a]. Overall, training evaluators is more prevalent for evaluating personalities and values due to the ease of associating textual expressions with specific personality and value dimensions.
More studies directly employ the LLM-as-a-judge approach [Gu et al., 2024] to score the open-ended responses. For example, Li et al. [2024d] prompt LLMs to score LLM responses across personality, ToM, and motivation. They validate the consistency between two LLM raters to indicate inter-rater reliability. Zheng et al. [2025] implement LLMs as personality evaluators and validate their consistency with human raters. The LLM-as-a-judge approach is particularly advantageous for evaluations in highly unstructured simulation settings. Zhou et al. [2024b], Wang et al. [2024d, a], Mou et al. [2024] apply LLM judges to assess SI across various dimensions, including goal completion, relationship maintenance, and adherence to social rules. Additionally, several works use embedding models to evaluate the similarity between LLM responses and prototypical examples [Huang et al., 2024c, Cahyawijaya et al., 2024, Amirizaniani et al., 2024].
@@ 6.5 Inference Parameters
Evaluation results and validation methods also depend on inference parameters of LLMs. Some studies use greedy decoding for deterministic outputs, requiring only a single response per item but sacrificing output diversity. In contrast, sampling-based decoding introduces stochasticity, generating a broader range of outputs by adjusting hyperparameters such as temperature, top-k, and top-p.
Inference parameters can affect the measured traits and abilities of LLMs. Diversifying the outputs may reveal a broader spectrum of latent traits, opinions, or cognitive strategies, but also introduce variability that complicates reliability and validity assessments. Conversely, deterministic settings enhance reproducibility but may obscure the model’s full range of capabilities or biases.
Most studies explicitly report and control these settings to ensure fair comparisons across models or experimental conditions. Some also examine the sensitivity of psychometric results to different decoding parameters. Transparent reporting and methodological rigor in parameter selection are essential. Investigating both deterministic and stochastic settings is recommended to fully understand the implications of test results.
@@ 7 Psychometric Validation
Figure 4: Overview of psychometric validation: reliability and consistency, validity, and standards and recommendations [Löhn et al., 2024].
Unlike AI benchmarking, which prioritizes system performance over test justification, psychometrics emphasizes theoretical grounding, standardized protocols, and reproducibility. Psychometric validation rigorously ensures the reliability, validity, and fairness of psychological tests. LLM psychometrics, as an emerging field, still lacks standardization for test design and administration, and recent research has begun exploring various aspects of psychometric validation.
@@ 7.1 Reliability and Consistency
Reliability is a fundamental principle in psychometric validation, assessing the degree to which a test is free from error. It encompasses consistency over time (test-retest reliability), across different versions (parallel forms reliability), and among evaluators (inter-rater reliability). Researchers have adapted these psychometric reliability metrics for LLM psychometrics. For instance, Li et al. [2024d] introduce a multi-dimensional psychometrics benchmark that extends to five forms of reliability, addressing the unique challenges of LLM evaluation: internal consistency, parallel forms reliability, inter-rater reliability, option position robustness, and adversarial attack robustness. Studies by Zheng et al. [2025], Huang et al. [2023a], Ceron et al. [2024], Shu et al. [2024] systematically investigate the consistency of LLM outputs under repeated trials, variations in prompt order, and cross-lingual contexts. They employ reliability indices such as Cronbach’s alpha, intraclass correlation coefficients, and agreement statistics to quantify both internal and external reliability dimensions. We note that some metrics pertain to both stochastic and systematic errors in LLM outputs; certain inconsistencies may compromise both the reliability and validity of LLM psychometrics.
Promising results indicate that some advanced LLMs consistently generate stable responses on the Big Five personality traits across diverse settings [Huang et al., 2023a]. Larger and instruction fine-tuned models exhibit even higher reliability when utilizing role-playing prompts [Serapio-García et al., 2023]. Moore et al. [2024] also observe relatively stable value orientations of LLMs across (1) paraphrases of a single question, (2) related questions within one topic, (3) multiple-choice and open-ended formats of a single question, and (4) multilingual translations. Li et al. [2024d] report strong inter-rater reliability when using LLMs as judges, though they note varying degrees of personality stability across models and differing levels of value consistency across cultural contexts.
@@
However, other studies present less favorable results. LLMs are notorious for prompt sensitivity in the context of psychometric evaluation. Trivial changes to prompt form [Röttger et al., 2024, Ren et al., 2024], option order [Li et al., 2024d, Dominguez-Olmedo et al., 2024], or syntactic structure [Hu and Levy, 2023] frequently result in systematic shifts in LLM outputs, severely undermining parallel forms reliability. Additionally, studies reveal distributional inconsistencies between next-token logits and forced-choice responses in psycholinguistic tests [Hu and Levy, 2023], as well as discrepancies between forced-choice and free-form responses in value [Kovač et al., 2024, Ye et al., 2025a] and political opinion surveys [Röttger et al., 2024]. In cognitive tests, Ullman [2023], Shapira et al. [2023a], Holterman and Deemter [2023] demonstrate that minor perturbations in ToM tasks often result in poor performance, indicating low parallel forms reliability of some ToM tests and LLMs’ dependence on fragile heuristics. Xu et al. [2025b] develop the Words and Deeds Consistency Test (WDCT) to measure the consistency between LLMs’ verbal and behavioral responses across various domains, including opinion vs. action, non-ethical value vs. action, ethical value vs. action, and theory vs. application. Their findings reveal that LLMs’ responses are inconsistent.
@@
The findings regarding reliability are clearly mixed. It seems to be influenced by various factors such as the target construct, forms of reliability, test formats, prompting, and the models used. LLMs tend to yield more reliable results in personality assessments [Huang et al., 2023a, Serapio-García et al., 2023] compared to political opinion evaluations [Dominguez-Olmedo et al., 2024, Röttger et al., 2024]. The degree of inter-rater agreement differs across various instruments and domains [Bodroža et al., 2024]. LLMs are reported to exhibit high internal consistency (Cronbach’s alpha $>0.8$), particularly in closed-choice personality tests [Huang et al., 2023a, Serapio-García et al., 2023, Zheng et al., 2025], while demonstrating low parallel form reliability due to prompt sensitivity [Li et al., 2024d, Shu et al., 2024, Gupta et al., 2024]. In addition, LLMs show greater consistency in structured tests with strong prompt control [Rozen et al., 2024, Serapio-García et al., 2023, Klinkert et al., 2024], but their reliability diminishes with less standardized prompts [Petrov et al., 2024] or when addressing more controversial topics [Moore et al., 2024]. The reliability also varies across models [Bodroža et al., 2024]. Moore et al. [2024] report that base models exhibit more consistent values compared to fine-tuned models in value-laden questions. Enhanced model safety is strongly associated with higher value consistency [Ye et al., 2025b], and more advanced models demonstrate superior role-playing performance, thereby increasing reliability in personality tests utilizing role-playing prompts [Serapio-García et al., 2023].

### v2

@@ 5.2 Data and task sources
Table 6: Overview of data and task sources: established inventories, custom-curated items, and synthetic items. For each data source, representative examples are provided for each construct.
@@
Established psychometric inventories offer standardized, well-validated tools for LLM evaluation. Common examples include BFI and HEXACO for personality [Serapio-García et al., 2023, Barua et al., 2024]; SVS, PVQ, WVS, and VSM for values [Fischer et al., 2023, Miotto et al., 2022]; MFT and DIT for morality [Abdulhai et al., 2024, Khandelwal et al., 2024]; ANES and PCT for political attitudes [Argyle et al., 2023, Bernardelle et al., 2024]; CRT for heuristics and biases [Hagendorff et al., 2023, Binz and Schulz, 2023]; and False-Belief Tasks for ToM [Ullman, 2023, Chen et al., 2024d]. While these inventories enable straightforward, standardized assessment, their use with LLMs raises concerns about data contamination, response bias, and limited ecological validity for complex AI tasks [Ye et al., 2025a].
@@
Human-authored, custom-curated items offer tailored psychometric tests that are often more relevant to LLMs, enabling exploration of novel capability dimensions. For example, Shen et al. [2024] find SVS and PVQ insufficient for LLM value alignment, so they create 11 AI-informed value statements based on a systematic review of alignment literature. Ceron et al. [2024] highlight the unreliability of standard political questionnaires for LLMs and introduce tests using annotated voting-advice questionnaires from seven EU countries to assess stance consistency. For morality, Hendrycks et al. [2021] curate a dataset of over 130,000 open-world scenarios requiring moral judgments. In ToM, Soubki et al. [2024] develop a dataset from real spoken dialogues to address gaps between synthetic benchmarks and human behavior. While custom-curated items offer high ecological validity, their development and validation are labor-intensive, limiting scalability and diversity.
@@
Synthetic items—generated primarily by LLMs—are an emerging approach in LLM Psychometrics, enabling large-scale, diverse, and contextually rich test content. This method requires careful prompt engineering and validation. Some studies adapt established inventories to improve ecological validity, reduce data contamination, and enhance scalability. For instance, Ren et al. [2024] convert self-report items into advice-seeking queries to better reflect real-world human-AI interactions. Bhandari et al. [2025b] rewrite established items into semantically equivalent forms, validating them via semantic embedding similarity thresholds. Zhu et al. [2024a] diversify cognitive test items to enable more comprehensive analysis of LLM cognitive abilities. Other work generates synthetic tests entirely from scratch. Hadar-Shoval et al. [2025] demonstrate that LLM-generated cognitive tests correlate strongly with human performance on established tests. Ye et al. [2025a] create value-eliciting prompts and confirm the reliability and validity of the resulting measurements. Jiang et al. [2024a] introduce generative evolving testing, where an LLM-based Item Generator produces items of specified difficulty. Chiu et al. [2025] use GPT-4 to generate moral dilemmas for evaluating LLM moral judgment. For social intelligence, agentic simulations prompt LLMs to generate components of social interactions [Zhou et al., 2024b, Mou et al., 2024].
@@ 5.3 Prompting strategies
Role-playing prompts (persona or profiling prompts) embed demographic or personal attributes into the LLM context during testing. These prompts enable the generation of multiple simulated participants from a single LLM for statistical analysis. For example, Serapio-García et al. [2023] prepend persona instructions to each item to elicit diverse responses, allowing examining reliability and validity within one LLM. Similarly, Ye et al. [2025b] use value-anchoring prompts to create hundreds of LLM-based participants for value measurement, supporting the development of an LLM-specific value system. LLMs can represent a superposition of personalities, values, and perspectives [Kovač et al., 2023], prompting research into their adaptability. Role-playing prompts show that LLMs can shift between personality types, though steerability varies by model and dimension [Lu et al., 2023, Jiang et al., 2023, La Cava and Tagarelli, 2024, Li et al., 2024d]. Such prompts are also used to test the stability of LLM values [Kovač et al., 2024, Rozen et al., 2024] and to measure how well LLMs can mimic moral or political profiles and their susceptibility to related biases [Münker, 2024, Simmons, 2023, Wright et al., 2024]. Cognitive benchmarks, especially for social interactions, often use persona prompts in agentic simulations [Zhou et al., 2024b, Mou et al., 2024, Huang et al., 2023b], with task performance depending on the LLM’s ability to embody these roles. These prompts influence social-cognitive reasoning when controlling for other variables [Tan et al., 2024]. While role-playing capabilities enable diverse perspectives useful for social simulations, they can also lead to unstable values and perspectives, potentially compromising test validity and complicating value alignment.
@@
Performance-enhancing prompts are crafted to augment LLMs in psychometric evaluations, as recommended by Hagendorff [2023]. Chain of Thought (CoT) prompting [Wei et al., 2022] reduces biases [Hagendorff et al., 2023], enhances social intelligence [Shapira et al., 2023a], and improves cognitive capabilities [Coda-Forno et al., 2024]. Its variant Emotional CoT [Li et al., 2024e] boosts EI, while emotional prompts [Li et al., 2023a] enhance general cognitive abilities through emotional stimuli. Few-shot prompting [Brown et al., 2020] has demonstrated improvements in ToM performance [Moghaddam and Honey, 2023]. Research has also developed construct-specific prompting strategies. Zhou et al. [2024a] propose MFT-guided reasoning prompts for moral alignment, while Echterhoff et al. [2024], Sumita et al. [2024] focus on self-debiasing approaches. ToM-specific enhancements include SymbolicToM [Sclar et al., 2023] and temporal decomposition-based prompting [Sarangi et al., 2025, Hou et al., 2024]. Additionally, Zhao et al. [2025a] develop a framework using self-reflection prompts to examine both explicit and implicit social biases, where explicit bias measurement reflects implicit bias.
@@ 5.4.1 Closed-ended output and scoring
LLM performance on structured tests is typically evaluated using logit-based probabilistic analysis and closed-ended output scoring. In multiple-choice or Likert-scale settings, some studies analyze token-level logits—especially the first token—to infer latent traits, detect response entropy, or compare LLM output distributions against human data [Pellert et al., 2024, Santurkar et al., 2023, Dominguez-Olmedo et al., 2024, Arehalli et al., 2022]. Psycholinguistic tests often use logit-based metrics such as surprisal (the negative log probability of a token), which quantifies predictability and cognitive effort, allowing researchers to align LLM uncertainty patterns with those of humans [Wang et al., 2024b, Steuer et al., 2023].
@@ 5.4.2 Open-ended output and scoring
Scoring open-ended outputs is more complex and typically falls into three categories: 1) rule-based, 2) model-based, and 3) human scoring.
@@
Rule-based scoring predominantly relies on the lexical hypothesis [Allport and Odbert, 1936], analyzing the presence and frequency of specific keywords or phrases. For example, Jiang et al. [2024b] apply LIWC (Linguistic Inquiry and Word Count) features [Pennebaker et al., 2001] to evaluate LLM personalities, and Fischer et al. [2023] employ a theory-driven value dictionary [Ponizovskiy et al., 2020] for value measurement. However, such lexicon-based approaches struggle to capture semantic nuance [Ye et al., 2025a]. Some rule-based methods go beyond keyword matching; for instance, Healey et al. [2024] detect bias by checking for deviations from equivalent treatment in responses.
@@
Model-based scoring is prevalent in unstructured testing for its flexibility and scalability. Many studies train models to evaluate LLM responses, such as fine-tuning BERT variants on the MyPersonality dataset for personality scoring [Hilliard et al., 2024]. Others fine-tune LLMs with psychometric inventories or human annotations to classify value valence [Ye et al., 2025a, Yao et al., 2024, 2025b, Sorensen et al., 2024a]. Scoring can be conducted at the item level (e.g., Generative Psychometrics: item parsing, scoring, aggregation [Ye et al., 2025a]) or at the response level, evaluating the entire output [Yao et al., 2024, 2025b]. Some models are tailored to specific frameworks (e.g., Schwartz’s Value Theory [Yao et al., 2024]), while others are trained to be generalists by leveraging broader fine-tuning data and LLM prior knowledge [Ye et al., 2025a]. Model-based evaluators are especially common for personality and value assessment, where textual cues map readily to target dimensions. Many studies now use the LLM-as-a-judge approach [Gu et al., 2024] for scoring open-ended responses. For instance, Li et al. [2024d], Zheng et al. [2025] prompt LLMs to rate responses on personality, ToM, and motivation, validating consistency with other LLMs or human raters. This method is especially useful in unstructured simulation settings, such as SI measurement across goal completion, relationship maintenance, and social rule adherence [Zhou et al., 2024b, Wang et al., 2024d, a, Mou et al., 2024]. Embedding models are also used to measure similarity between LLM outputs and prototypical examples [Huang et al., 2024c, Cahyawijaya et al., 2024, Amirizaniani et al., 2024].
@@ 5.5 Inference parameters
Evaluation and validation depend on LLM inference parameters. Greedy decoding yields deterministic, reproducible outputs but limits diversity, while sampling-based decoding (adjusting temperature, top-k, top-p) introduces stochasticity and diversifies responses. Sampling-based decoding may reveal a broader spectrum of latent traits, opinions, or cognitive strategies, but also introduce variability that complicates psychometric validation. Conversely, deterministic settings enhance reproducibility but may obscure the model’s full range of capabilities or biases. Most studies report and control inference settings for fair comparisons, and some analyze sensitivity to these parameters. Transparent reporting and careful parameter selection are essential; both deterministic and stochastic settings should be considered to fully understand the implications of test results.
@@ 6 Psychometric validation
Figure 4: Overview of psychometric validation: reliability and consistency, validity, and standards and recommendations.
Unlike AI benchmarking, which focuses on system performance, psychometrics prioritizes theoretical grounding, standardized protocols, and reproducibility. Psychometric validation ensures that tests are reliable, valid, and fair. As an emerging field, LLM Psychometrics lacks standardization in test design and administration, and recent work is beginning to address these issues. We outline the landscape of psychometric validation in Fig. 4.
@@ 6.1 Reliability and consistency
Reliability, a core aspect of psychometric validation, measures how consistently a test performs—over time (test-retest), across versions (parallel forms), and among evaluators (inter-rater). Researchers have adapted these metrics for LLMs. For example, Li et al. [2024d] propose a benchmark covering five reliability forms: internal consistency, parallel forms, inter-rater, option position robustness, and adversarial attack robustness. Other studies [Zheng et al., 2025, Huang et al., 2023a, Ceron et al., 2024, Shu et al., 2024] examine LLM output consistency across repeated trials, prompt variations, and languages. Notably, some metrics capture both random and systematic errors, so inconsistencies can undermine both reliability and validity.
@@
Some advanced LLMs show stable Big Five personality trait responses across settings [Huang et al., 2023a], with larger and instruction-tuned models achieving even higher reliability using role-playing prompts [Serapio-García et al., 2023]. LLMs also display consistent value orientations across paraphrased, related, and translated questions, as well as different response formats [Moore et al., 2024]. Strong inter-rater reliability is reported when LLMs act as judges, though personality stability and value consistency vary by model and cultural context [Li et al., 2024d]. However, other studies report less favorable results. LLMs are notoriously sensitive to prompt variations: small changes in prompt form [Röttger et al., 2024, Ren et al., 2024], option order [Li et al., 2024d, Dominguez-Olmedo et al., 2024], or syntax [Hu and Levy, 2023] can systematically shift outputs, undermining parallel forms reliability. Studies also find inconsistencies between next-token logits and forced-choice responses in psycholinguistic tests [Hu and Levy, 2023], and between forced-choice and free-form responses in value [Kovač et al., 2024, Ye et al., 2025a] and political opinion surveys [Röttger et al., 2024]. In cognitive tests, minor changes to ToM tasks often cause poor performance, indicating low parallel forms reliability and reliance on fragile heuristics [Ullman, 2023, Shapira et al., 2023a, Holterman and Deemter, 2023]. The Words and Deeds Consistency Test (WDCT) [Xu et al., 2025b] further shows LLMs’ verbal and behavioral responses are inconsistent across domains.
@@
Reliability findings are mixed and depend on factors such as construct, reliability type, test format, prompting, and model. LLMs are generally more reliable in personality assessments [Huang et al., 2023a, Serapio-García et al., 2023] than in political opinion evaluations [Dominguez-Olmedo et al., 2024, Röttger et al., 2024]. Inter-rater agreement varies across instruments and domains [Bodroža et al., 2024]. LLMs often show high internal consistency (Cronbach’s alpha $>0.8$) in closed-choice personality tests [Huang et al., 2023a, Serapio-García et al., 2023, Zheng et al., 2025], but low parallel form reliability due to prompt sensitivity [Li et al., 2024d, Shu et al., 2024, Gupta et al., 2024]. Consistency improves in structured tests with strong prompt control [Rozen et al., 2024, Serapio-García et al., 2023, Klinkert et al., 2024], but declines with less standardized prompts [Petrov et al., 2024] or controversial topics [Moore et al., 2024]. Reliability also varies by model: base models are more consistent than fine-tuned ones on value-laden questions [Moore et al., 2024], and higher model safety correlates with greater value consistency [Ye et al., 2025b]. Advanced models also perform better in role-playing personality tests, increasing reliability [Serapio-García et al., 2023].